OpenStudy (mind2000):

4 is what percent of 5

OpenStudy (austinl):

4/5=8/10=.80=80%

ganeshie8 (ganeshie8):

another trick is to multiply both wid 20 :- 4 is what percent of 5 80 is what percent of 100

OpenStudy (austinl):

That is also a good trick, but I love ratios for some reason! They are just fun!

ganeshie8 (ganeshie8):

I net they're fun :)

ganeshie8 (ganeshie8):

*bet